Kevin is 16 years old and Vanessa is 2 years old.
How many years will it take until Kevin is only 3 times as old as Vanessa?

Answer by TimothyLamb(4379)

x = years in future when kevin is 3 times vanessa's age
---
16 + x = 3(2 + x)
16 + x = 6 + 3x
2x = 10
x = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php